OpenStudy (anonymous):

What is the expression written using each base only once? 5^-13 x 5^-5 x 2^2

OpenStudy (hba):

what answer did u get

OpenStudy (anonymous):

I didn't get an answer yet. Normally it would be 5^-8 if it was just the first two variables, but the extra 2^2 throws me off.

OpenStudy (anonymous):

Can anyone help me with this question?

OpenStudy (hba):

ok r u allowed to use the calculator

OpenStudy (anonymous):

No. Even if I could I couldn't rewrite the expression into another one using a calculator.

OpenStudy (lgbasallote):

2^2 means nothing...it's like x^3 and y^2

OpenStudy (lgbasallote):

\[\Large x^3 \cdot y^2 \implies x^3y^2\]

OpenStudy (lgbasallote):

btw -8 is wrong

OpenStudy (lgbasallote):

\[\large 5^{-13} \times 5^{-5} \implies 5^{-13 + -5}\]

OpenStudy (anonymous):

\[\large 5^{-13}\times 5^{-5} \times 2^2\] \[\large 5^{-13-5}\times 2^2\]

OpenStudy (anonymous):

So it would be 5^-17 and then multiply that with 2^2, which would the final value would be 10^-15 or what?

OpenStudy (lgbasallote):

still not -17

OpenStudy (lgbasallote):

what is 13 + 5?

OpenStudy (anonymous):

Oh 18 my bad haha

OpenStudy (anonymous):

also 5^(-17) can't be multiplied to 2^2 like that. you have to do them seperately.

OpenStudy (lgbasallote):

so what's -13 - 5?

OpenStudy (anonymous):

-18. And to phuong, so I leave the 2^2 separately?

OpenStudy (lgbasallote):

like i said before you cant combine x^2 and y^3

OpenStudy (lgbasallote):

you just put them side-by-side

OpenStudy (lgbasallote):

i.e \[4^2 \times 3^3 = 4^2 3^3\]

OpenStudy (lgbasallote):

unless you have a calculator then you input those numbers

OpenStudy (anonymous):

Ah I see. So (5^18)(2^2) would be the answer?

OpenStudy (lgbasallote):

nope not 18

OpenStudy (lgbasallote):

missing soemthing ;)

OpenStudy (anonymous):

Ah I'm sorry I keep forgetting! (5^-18)

OpenStudy (lgbasallote):

better

OpenStudy (anonymous):

Yeah thanks man. I appreciate your help guys!